OpenStudy (anonymous):

During one week in the winter, the temperature dropped an average of 2° per hour for 8 hours. If the temperature was -2°F at the beginning of the 8 hours, what was the temperature at the end?

OpenStudy (anonymous):

Come on man, at least have a shot at figuring it out :-P -2 + (8*-2) = -2 - 16 = -18